OpenStudy (anonymous):

suppose you have a bag that holds 20 pieces of candy. 10 grapes, 2 cream soda, and 8 apple. you reach into the bag pull out one piece. find the probability the the piece you pull out is apple flavored

OpenStudy (anonymous):

there are 8 apples out of 20 candies so P(apple) = 8/20 or 2/5

OpenStudy (anonymous):

suppose you roll 2 dice, find the probability of not rolling a sum of 12

OpenStudy (anonymous):

probability of 12 (2 6's) is 1/6 * 1/6 = 1/36 P (not 12) = 1 - 1/36 = 35/36